OpenStudy (anonymous):

Use synthetic division to find P(3) for P(x) = x^4 – 6x^3 – 4x^2 – 6x – 2. Multiple Choice Answers 3 –137 299 –47

OpenStudy (anonymous):

do you know how to do synthetic div?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

get you calculator check you answer P(x) = x^4 – 6x^3 – 4x^2 – 6x – 2. let x=3 P(3) = 3^4 – 6(3)^3 – 4(3)^2 – 6(3) – 2=___?. alternate use synthetic div 3 [ +1 -6 - 4 -6 -2 +3 -9 -39 -135 ------------------------------------ 1 -3 -13 -45 R=-137

OpenStudy (anonymous):

look them here also https://www.google.com.ph/search?q=synthetic+division&hl=fil&client=firefox- beta&hs=ZaJ&rls=org.mozilla:en-US:official&prmd=imvns&tbm=isch&tbo=u&source=univ&sa=X&ei=GGWiUPv0DebKmgWIuIDoCQ&ved=0CDoQsAQ&biw=1024&bih=605&sei=SWWiUK-sFsWcmQW-24DQBA

OpenStudy (anonymous):

my internet is almost gone i got to go now ...good luck now...:D